A window washer with a mass of 67.0 kg stands a distance, D = 0.800 m, from the left end of a plank of length, L= 2.10 m, with a mass of 21.0 kg. The plank is hung on two cables. Find T2, the tension in the right cable.
Answer #1

  
Tension in each cable, no window washer = (21 x 9.8)/2, = 102.9N.
Now the window washer, 89kg. at .4m. from the left end.
(67 x 9.8) = 656.6N.
(2.1m/0.8m) = 2.62:1 ratio. Add both sides together, = 3.63
(656.6/3.63) = 180.881+ 102.9 = 283.78N. tension in right cable.